Show that if x is a real number, then x − 1 < ⌊x⌋ ≤ x ≤



⌈x⌉ < x + 1.

1
Expert's answer
2021-11-29T05:34:49-0500

"n=\lfloor x\rfloor\quad \Leftrightarrow \quad n\leq x<n+1
\quad \Leftrightarrow \quad \begin{cases}
x-1<n
\\
n\leq x
\end{cases}
\quad \Leftrightarrow \quad 
x-1<
\lfloor x\rfloor \leq x"

"m=\lceil x\rceil\quad ⇔\quad m-1<x\leq m \quad \Leftrightarrow \quad \begin{cases}
x\leq m
\\
m<x+1
\end{cases}
\quad \Leftrightarrow \quad 
x\leq 
\lceil x\rceil <x+1"


So, "x-1<\lfloor x \rfloor \leq x\leq \lceil x \rceil <x+1"